OpenStudy (anonymous):

what is the anti derivative of f (x) = 12x^3 -9x^2 + 14x - 4 ?

OpenStudy (sburchette):

\[\int\limits(12x^3-9x^2+14x-4)dx\] You just have to integrate each term using the power rule for integration which is\[\int\limits(cx^n)dx=(c/(n+1))x^{n+1}+C\] Where c and n are real constants. Using the rule we get \[3x^4-3x^3+7x^2-4x+C\]

OpenStudy (anonymous):

thats what i thought thanks for confirming!!

OpenStudy (sburchette):

My pleasure =)